# arXiv:2603.03220v1[cond-mat.mtrl-sci] 03 Mar 2026

### v1

@@ 3.1 Korringa-Kohn-Rostoker Green’s function method
We calculate the Korringa-Kohn-Rostoker (KKR) Green’s function (GF) for a NC ground state following the standard approach [Mankovsky2011, Simon2019]. First the single site scattering problem is solved yielding the corresponding regular and irregular solutions, $R_{\sigma L\sigma^{\prime}L^{\prime}}^{m}(r_{m},z)$ and $H_{\sigma L\sigma^{\prime}L^{\prime}}^{m}(r_{m},z)$ as well as the corresponding $t$-matrix. Following this, the backscattering operator taking into account the multiple electron-scattering in the lattice is found as
@@
Einstein sum convention is used in equations 25 and 26. The KKR approach yields an all-electron picture of the Kohn-Sham system with the core states and atomic potentials found self-consistently. We assume the system to feature the discrete translational symmetry including the magnetization orientation as well and, thus, we do not deploy the generalized Bloch theorem for incommensurate spirals in the sense of Sandratskii [Sandratskii1991].
@@ 3.2 Evaluation of susceptibility
Here, ${\mathbf{r}}_{s}$ should be understood as a reduced vector $\mathbf{r}$}, ${\mathbf{r}}_{s}\equiv{\mathbf{r}}-{\mathbf{s}}_{s}$, where ${\mathbf{s}}_{s}$ is the center of the atomic polyhedron containing $\mathbf{r}$}. The angular dependence of the spatial arguments in each cell is developed using spherical harmonics ${\mathcal{Y}}_{L}$ with index $L\equiv lm$ and argument $\hat{{\mathbf{r}}}_{s}\equiv{\mathbf{r}}_{s}/{\left|{\mathbf{r}}_{s}\right|}$. Note that in the case of full potential calculations, the product of two shape functions for the polyhedra $s$ and $s^{\prime}$ is incorporated into the definition of $\chi^{ss^{\prime}}_{LL^{\prime}}{\left(r_{s},r_{s^{\prime}}\right)}$. The radial dependence of $\chi{\left(r_{s},r_{s^{\prime}}\right)}$ can be developed in different ways. In our previous works, cf., e.g., [Buczek2011a], we took advantage of Chebyshev polynomials in order to achieve a rapid convergence with respect to the number of basis functions, as, typically, only around 12 were sufficient to fairly describe the dependence. In the current work, the function $\chi{\left(r_{s},r_{s^{\prime}}\right)}$ is given directly in the real space on a Gaussian mesh between 0 and the radius of the bounding atomic sphere $s$. This approach leads to somewhat increased memory consumption but greatly enhances the fulfilment of the Goldstone sum rule. Upon this step, the susceptibilities are represented as ordinary two-dimensional matrices $\chi_{\lambda\lambda^{\prime}}$ where the super-index $\lambda={\left(i,s,L,\rho\right)}$ comprises the density channel designation $i$, site $s$, angular quantum numbers $L\equiv lm$, and the radial index $\rho$. In this representation, Eq. (2.2) can be solved by the following direct matrix inversion [Staunton1999, Staunton2000, Thakor2003].
@@
4 Magnons of IrMn3
@@
Let us apply our formalism to the TAF IrMn\$_3\$, in which the Mn atoms, exhibiting a finite magnetic moment, form a NC structure. More precisely, we take a look at the characteristic 111 plane of the system. Figure 1 shows the magnon dispersion and the inverse lifetimes (shown as bars designating the full with at half-maximum, FWHM, of the magnon peak) obtained from our LRTDDFT calculations. Three linearly dispersing magnon modes emerge in the long wave-length limit from three distinct Goldstone modes, as expected from the symmetry analysis outlined above. Their maximum energies at the edge of the Brillouin zone amounts to about 370meV. The modes are clearly Landau damped but remain well-defined for all momenta. In general, the damping increases with energy but differs strongly depending on the magnon polarization.
Figure 1: Magnon-spectrum within the 2D Brillouin zone, i.e. magnons propagating in the 111-lattice plane. Error bars inform about the FWHM of the magnon peak, i.e. the inverse lifetime. On the right, the 2D Fermi surface is shown together with the path through the Brillouin zone.
5 Summary and discussion
@@
We presented a novel implementation of the ab initio linear response time-dependent density functional theory addressing spin excitations in non-collinear magnets based on the Korringa-Kohn-Rostoker Green’s function method. We elaborated the details of the numerical implementation, focusing on the managing the algorithmic complexity by means of symbolic computational algebra, exposed the generalization of the adiabatic local spin density approximation to the exchange-correlation kernel to the non-collinear case. The emergence of multiple Nambu-Goldstone bosons in non-collinear susceptibility calculations is elaborated on formally and from the numerical convergence point of view, linked to the zero-energy magnetic deformation eigenvectors. With the aid of the formalism, we analysed the dispersion and Landau damping of magnons in kagome TAF Mn3Ir. We hope that this work will spark further developments in the non-collinear magnonics.
@@
Acknowledgments. We thank Leonid M. Sandratskii, László Szunyogh, Patrick Perndorfer, and Sebastian Paischer for fruitful discussions and Igor Maznichenko for his advice and helpful remarks. P.B. and A.E. acknowledge the funding, respectively, by Österreichischer Fonds zur Förderung der Wissenschaftlichen Forschung (FWF) under grant I 5384 and by Deutsche Forschungsgemeinschaft (DFG) under grant BU 4062/1-1.

### v2

@@ 3.1 Korringa-Kohn-Rostoker Green’s function method
The generalization of the Korringa-Kohn-Rostoker (KKR) Green’s function (GF) method to NC ground states is based on the following approach [Mankovsky2011, Simon2019]. First, the single site scattering problem is solved yielding the corresponding regular and irregular solutions, $R_{\sigma L\sigma^{\prime}L^{\prime}}^{m}(r_{m},z)$ and $H_{\sigma L\sigma^{\prime}L^{\prime}}^{m}(r_{m},z)$ as well as the corresponding $t$-matrix. Subsequently, the backscattering operator taking into account the multiple electron-scattering in the lattice is found as
@@
Einstein sum convention is used in equations (25) and (26). The KKR approach yields an all-electron picture of the Kohn-Sham system with the core states and atomic potentials found self-consistently. We assume the system to feature the discrete translational symmetry including the magnetization orientation, thus, we do not have to deploy the generalized Bloch theorem for incommensurate spirals in the spirit of Sandratskii [Sandratskii1991].
@@ 3.2 Evaluation of susceptibility
Here, ${\mathbf{r}}_{s}$ should be understood as a reduced vector $\mathbf{r}$}, ${\mathbf{r}}_{s}\equiv{\mathbf{r}}-{\mathbf{s}}_{s}$, where ${\mathbf{s}}_{s}$ is the center of the atomic polyhedron containing $\mathbf{r}$}. The angular dependence of the spatial arguments in each cell is developed using spherical harmonics ${\mathcal{Y}}_{L}$ with index $L\equiv lm$ and argument $\hat{{\mathbf{r}}}_{s}\equiv{\mathbf{r}}_{s}/{\left|{\mathbf{r}}_{s}\right|}$. Note that in the case of full potential calculations, the product of two shape functions for the polyhedra $s$ and $s^{\prime}$ is incorporated into the definition of $\chi^{ss^{\prime}}_{LL^{\prime}}{\left(r_{s},r_{s^{\prime}}\right)}$. The radial dependence of $\chi{\left(r_{s},r_{s^{\prime}}\right)}$ can be developed in different ways. In our previous works, cf., e.g., [Buczek2011a], we took advantage of Chebyshev polynomials in order to achieve a rapid convergence with respect to the number of basis functions, as, typically, only around 12 were sufficient to fairly describe the dependence. In the current work, the function $\chi{\left(r_{s},r_{s^{\prime}}\right)}$ is given directly in the real space on a Gaussian mesh between 0 and the radius of the bounding atomic sphere $s$. This approach leads to somewhat increased memory consumption but greatly enhances the fulfillment of the Goldstone sum rule. Upon this step, the susceptibilities are represented as ordinary two-dimensional matrices $\chi_{\lambda\lambda^{\prime}}$ where the super-index $\lambda={\left(i,s,L,\rho\right)}$ comprises the density channel designation $i$, site $s$, angular quantum numbers $L\equiv lm$, and the radial index $\rho$. In this representation, Eq. (2.2) can be solved by the following direct matrix inversion [Staunton1999, Staunton2000, Thakor2003]:
@@
3.3 Antiferromagnetic magnon peak fitting
As pointed out by Skovhus and Olsen [Skovhus2022], the magnon peak fitting in the case of the antiferromagnets is more challenging than for the ferromagnets, especially for small momenta. It is so due to the characteristic presence of the two peaks anti-symmetric with respect to $\omega=0$ becoming very close to each other in the short-wavelength regime, see Fig. 1. Principally, we analyze and fit the magnon peaks on the nearly real axis, i.e., along a line on the upper complex semi-plane parallel to the real axis and in the distance $\gamma$ from it. The susceptibility is directly evaluated for points $\omega+{{\mathfrak{i}}}\gamma$ and no intermediate analytic continuation to real real axis ($\omega+{{\mathfrak{i}}}{0^{+}}$) is performed. This is for two reasons. First, the analytic continuation is known to be numerically ill-conditioned and unstable. Despite extended research, we identified no reliable, robust, and automatable numerical algorithm for its execution. Second, weakly damped magnons appear as poles very close to the real axis with corresponding significant eigenvalue magnitude and extremely narrow Lorenz-like peak practically precluding correct fit. In this work, referring to [Skovhus2022], we deploy the following magnon line ansatz
$a_{\lambda}{\left({\mathbf{q}},\omega\right)}=-A_{\lambda{\mathbf{q}}}{\left(\frac{1}{{\left(\omega-\omega_{\lambda{\mathbf{q}}}\right)}^{2}+{\left(\gamma+\eta_{\lambda{\mathbf{q}}}\right)}^{2}}-\frac{1}{{\left(\omega+\omega_{\lambda{\mathbf{q}}}\right)}^{2}+{\left(\gamma+\eta_{\lambda{\mathbf{q}}}\right)}^{2}}\right)}.$ (32)
Here, $A_{\lambda{\mathbf{q}}}$ stands for the spectral weight of the magnon mode, $\omega_{\lambda{\mathbf{q}}}$ is the magnon frequency, and $\eta_{\lambda{\mathbf{q}}}$ the intrinsic magnon decay parameter. We recall $\gamma$ to be the distance of the nearly real axis, i.e., the artificial broadening. For large $\omega_{\lambda{\mathbf{q}}}$, the peak’s maximum appears approximately at this frequency, and the full width at half-maximum (FWHM) of the peak amounts to roughly $2\eta_{\lambda{\mathbf{q}}}$. As carefully discussed in [Skovhus2022], for $\omega_{\lambda{\mathbf{q}}}\to 0$, the peak’s maximum and FWHM do not directly yield these fit parameters.
On the nearly real axis, the eigenvalues of the $L(\omega+{{\mathfrak{i}}}\gamma)$ do not necessarily vanish for $\omega=0$, contrary to the requirement of Eq. (32). This difficulty can be circumvented upon considering the sum of eigenvalues corresponding to each other for positive and negative frequencies. As evident from Fig. 1, the sum goes through the origin and is represented well by the equation.
Although the ansatz in Eq. (32) includes only three parameters, the quality of the fit and the stability of the results for varying $\mathbf{q}$}, especially in the long wavelength regime, improve considerably with the number of points in the fitting set. The low frequency region (shaded blue region in Fig. 1) turns out to be particularly relevant and in our calculations we typically increase the frequency point density there.
Figure 1: Magnon peak fitting in the case of enhanced susceptibility. $\times$, ++, and $\cdot$ points correspond to the three relevant, i.e., of significant value, magnon mode eigenvalues of the loss matrix evaluated in our LRTDDFT scheme for energies on the nearly real axis. Following the properties of the Fourier-transformed response function, cf. Appendix  A, the magnon peaks show typical anti-symmetry with respect to $\omega=0$. In the blue region, the energy mesh density is increased to stabilize the peak parameters’ fit. Note that the mesh density shown in the plot, for better visibility, is only a fraction of the actual density.
3.4 Convergence considerations
Figure 2: Convergence of the enhanced susceptibility with respect to the number of $\mathbf{k}$} points used for the Brillouin zone integration. For the case of Mn3Ir, the most dense mesh necessary has $N_{k}\times N_{k}\times N_{k}$ points with $N_{k}=25$. The three groups of lines correspond to the three relevant magnon eigenvalues.
Let us conclude this section with several remarks regarding the numerical convergence properties of our scheme. They resemble much the ones encountered in the collinear case [Buczek2011a]. The most computationally expensive part of the algorithm is the Brillouin zone integration of the product of two backscattering operators, i.e., their $\mathbf{k}$}-convolution. For the case of Mn3Ir, the most dense mesh necessary has $N_{k}\times N_{k}\times N_{k}$ points with $N_{k}=25$, cf. 2. Note that this number is reduced considerably once the backscattering operators are evaluated away from the real axis. Furthermore, the number $N_{k}=25$ is moderate compared to other systems. For example, the bcc Fe requires $N_{k}=100$ or more. This is due to the fact, plainly speaking, that Fe features large density of states at the Fermi energy. The strongly dispersing bands of Mn3Ir crossing the Fermi level result in a flat and low density of states in this region and the relaxed convergence demand.
In the spherical harmonics expansion of the KKR Green’s function it is enough to include only states up to $d$ ($l_{\mathrm{max}}^{\mathrm{KKR}}=2$) when computing the density response functions. When concentrating on the dynamics of moments as a whole, the expansion of the response functions with $l_{\mathrm{max}}^{\mathrm{susc}}=0$ yields already accurate enough results. The cases of $l_{\mathrm{max}}^{\mathrm{susc}}=1$ and $l_{\mathrm{max}}^{\mathrm{susc}}=2$ are indiscernible and differ only weakly from the spherical approximation.
We find, however, that the radial dependence of susceptibility components in atomic cell influences the accuracy of the calculation, in particular the Goldstone divergence, strongly. The radial mesh points are chosen according to the Gauss-Legendre quadrature scheme. A Neville interpolator is used in order to retrieve the values of the radial solutions on the Gauss mesh from the denser logarithmic mesh which is used in the underlying KKR calculations. It turns out that around 20 points per atom balance accuracy and computational cost reasonably in the case of Mn3Ir.
4 Results
4.1 Electronic ground state
Figure 3: $111$-plane of Mn3Ir revealing the characteristic non-collinear KAFM magnetic pattern. The staking of these shifted planes generates the entire crystal. The color coding of the magnetic Mn atoms is consistently maintained throughout the paper. The global coordinate system is oriented such that the $z$-axis points out of the $111$-plane and $x$ and $y$ axes lie in this plane. The local coordinate systems for each Mn atom have their $z$-axes along the moments’ directions, the $y$-axes are assumed to lie in the kagome $111$-plane, and $x$-axes are chosen such that right-handed coordinate systems form. The Mn2 atomic magnetic moment is taken to be oriented in the global $y$ direction.
Fig. 3 depicts the kagome $111$-plane of Mn3Ir investigated in this paper. The lattice constant of the underlying cubit lattice is taken to be a=3.785a=3.785Å. The Mn-Mn atomic distance amounts to $a/\sqrt{2}$ and the distance between 111-planes to $a/\sqrt{3}$. The global coordinate system is oriented such that $z$-axis points out of the $111$-plane and $x$ and $y$ axes lie in this plane. The local coordinate systems for each Mn atom have their $z$-axes along the moments’ directions, the $y$-axes are assumed to lie in the kagome $111$-plane, and $x$-axes are chosen such that right-handed coordinate systems form. The Mn2 atomic magnetic moment is taken to be oriented in the global $y$ direction.
Figure 4: Electronic Bloch spectral function showing the contributions of the Mn atoms to the total band structure. The directions within the Brillouin zone can be found as an inset in Fig. 6. Color encodes information about the spin polarization (magnetization direction) of the Bloch state for a particular electronic momentum $\mathbf{k}$}. The intensity of the band signifies the Bloch amplitude of the state on the atom.
The resulting band structure is intrinsically spin-polarized despite its vanishing net magnetization [3, 4], cf. Fig. 4, and the altermagnetic symmetry of the band structure is clearly discernible. It originates from the fact that magnetic moments of different directions are embedded in different crystal environments imposed by the kagome pattern. The altermagnetic symmetry has one more consequence: The amplitude of the Bloch states (electronic occupation and spin polarization) on particular magnetic atoms can differ even for wave vectors related by the symmetry of the underlying non-magnetic lattice. This is clearly seen in Fig. 4 for Mn2 and Mn3 atoms as asymmetry between $K$ and $K^{\prime}$ directions. At $\Gamma$, one clearly identifies occupied electronic bands around 400meV below the Fermi level polarized along the respective moments’ directions. They are responsible for the moments’ formation. The Fermi surface is formed by strongly dispersing bands of heterogeneous spin polarizations. Above the Fermi level and away from the center of the Brillouin zone, the magnetic orientation of bands differs strongly from the moment direction. In our recent study [Eilmsteiner2026a], we have shown that this peculiarities of the NC altermagnetic band structure result in an intricate mode-selective Landau magnon damping.
4.2 Spectral analysis of density excitations
The spectrum of particle-hole and collective excitations is given by the loss matrix of, respectively, KS and enhanced susceptibility, cf. Eq. (3). Owing to the clear interpretation of the eigenvalues $a_{\lambda}{\left({\mathbf{q}},\omega\right)}$ of $L^{ij}{\left({\mathbf{r}},{\mathbf{r}}^{\prime},{\mathbf{q}},\omega\right)}$ we take them, in most cases, directly as the starting point for any subsequent analysis. Fig. 5 presents an example of the KS particle-hole spectrum for a selected momentum $\mathbf{q}$}. The series of data points for a given $\omega$ represents the eigenvalues $a_{\lambda}{\left({\mathbf{q}},\omega\right)}$ indexed by $\lambda$, each associated with a different eigenvector $\xi_{\lambda}^{i}{\left({\mathbf{r}},{\mathbf{q}},\omega\right)}$. For given $\left({\mathbf{q}},\omega\right)$}-pair, the set of $\left\{\xi_{\lambda}^{i}{\left({\mathbf{r}}\right)}\right\}$} forms an orthogonal and complete basis for representing charge and spin densities in the primitive cell of the crystal. In general, when arguments $\left({\mathbf{q}},\omega\right)$} change, a new set of eigenvectors is obtained. The index $\lambda$ runs between 1 and $\chi$-matrix order given by the choice of computational basis, roughly of order of $10^{3}$ for our kagome magnet. In Fig. 5, all eigenvalues with the magnitude over certain threshold are shown, the remaining being of minor weight. Still, we observe that the KS excitations, depicted in blue in figure 5a), are associated with a multitude of relevant eigenvectors (figure 5 b)) involving, in general, charge and magnetization components of different directions. The picture changes when the true (enhanced) response is considered, cf. red lines in  Fig. 5a), markers in  Fig. 5c), and the data in Fig. 1. In the energy window below 500meV, only a few eigenvalues of the enhanced susceptibility show a sizable magnitude, other being negligible. Fig. 5c) reveals that the corresponding modes practically involve only oscillating spin density oriented perpendicular to the local ground state magnetization direction, whereas the longitudinal and charge components exhibit vanishing weight. This is in good agreement with a basic assumption of the atomistic Heisenberg model, in which atomic magnetic moments are assumed to be of constant magnitude and charge degrees of freedom are neglected.
Figure 5: Eigenvalues (a) and some prominent eigenvectors (b,c) of the Loss matrix for a wave vector $\mathbf{q}$ in the middle of the Brillouin zone. The eigenvalues of the KS system shown in blue form a more flat spectrum while the eigenvalues of the interacting system, in red, exhibit a strong magnon peak. Analysis of the eigenvectors, given in the site local basis, shows the Heisenberg-like dynamical behavior of the interacting system, while the prominent eigenvectors of the KS system have non vanishing charge (0) and longitudinal (z) components.
In this vein, the magnons known from the Heisenberg model arise naturally in the dynamical susceptibility calculations. The three relevant eigenvectors correspond to the three magnon modes and will be discussed more carefully below. Interestingly, the eigenvalues for different frequencies $\omega$ can relatively easily be assigned to a particular mode because the magnon eigenvector $\xi_{\lambda}^{i}{\left({\mathbf{r}},{\mathbf{q}},\omega\right)}$ changes only weakly with the frequency. (However, we recall that for a given frequency $\omega$ the magnon eigenvectors $\xi_{\lambda}$ are orthogonal and thus much different.) We finally note that for energies above the magnon range, there additionally emerges a plethora of coupled charge and magnetization density excitations, e.g., longitudinal spin fluctuations and plasmons [Buczek2020]. They will be studied in a separate work.
4.3 Magnon dispersion and the chirality-selective Landau damping
@@
Figure 6 shows the magnon dispersion and the inverse lifetimes (shown as bars designating the full with at half-maximum, FWHM, of the magnon peak) obtained from our LRTDDFT calculations. Three linearly dispersing magnon modes emerge in the long wave-length limit from three distinct Goldstone modes, as expected from the symmetry analysis outlined above. Their spin wave velocities amount to around $v_{\Gamma}\approx 80$km/s and their maximum energies at the edge of the Brillouin zone reach about 370meV. The modes are clearly Landau damped but remain well-defined for all momenta and can be treated as underdamped harmonic oscillators. The maximal FWHM of a magnon mode is found to be about 140meV but most magnons live significantly longer. While the magnon dispersions of Mn3Ir and Mn3Rh are qualitatively similar, the magnons in the first system are of higher energies and stronger Landau damped than in the latter, cf. [Eilmsteiner2026a].
Figure 6: The spectrum of magnons in the kagome 111-plane of Mn3Ir. The color-coding of the magnon branches is used consistently in this section. (a) Magnon energies as functions of momentum. The tree branches are practically degenerate concerning the energies but not the mode-selective damping. Error bars signify the FWHM of the magnon peak, i.e. the inverse lifetime of the quasiparticle. In the inset, the 111-cut of the Fermi surface is shown together with the path through the Brillouin zone. (b) FWHM of the magnon peaks as a function of momentum. (c) FWHM of the magnon peaks as a function of the magnon energy. The mode-selective Landau attenuation is evident from Figs. b) and c).
Similarly to the case of Mn3Rh investigated recently [Eilmsteiner2026a], above the threshold of about 150meV, the FWHM of magnons peaks in Mn3Ir is a non-monotonous function of both momentum and magnon energy. This makes these two materials extremely promising candidates for the terahertz spintronic applications. The relatively low damped high-frequency modes, before their decay, perform a larger number of revolutions compared to their low-frequency companions, despite featuring larger attenuation parameter (FWHM). This feature will be further investigated in Sec. 4.4.
Another remarkable property of the Landau damping shared by the two kagome magnets is its pronounced mode selectivity. The decay rates differ significantly even for magnonic modes of practically the same energy and momentum. To be specific, let us consider the low-momentum magnons in Fig. 6b) Two modes ($\alpha$-type, green and yellow markers) are up to factor 3 weaker Landau damped than the third branch of $\beta$-type (magenta). As discussed carefully before [Eilmsteiner2026a], this selectivity can be traced to the spectrum of the non-collinear Stoner excitations resulting from the altermagnetic character of the underlying electronic band structure interplaying with the forms of the magnons’ eigenvectors.
4.4 Magnetic modes
Figure 7: Real-space and parametric real-time depiction of the magnetic moments’ precession associated with magnon eigenvectors obtained from our LRTDDFT calculations, including the Landau attenuation. The modes become excited at $t=0$. The three magnetic Mn atoms of the primitive cell are considered. The color coding corresponds to the magnon branches identified in the dispersion given in Fig. 6 Stars signify the starting orientation at $t=0$. For $t>0$, the components of the moments’ deviations in the respective local moment coordinate systems are plotted.
Figure 8: The same patterns of precession as in Fig. 7 but shown as explicit function of time. The solid lines stand for the local $x$ direction (out of 111-plane) while the dashed ones for the local $y$ direction (in 111-plane).
Let us discuss the real-space and real-time dynamics of magnons modes identified in the preceding subsection. In our analysis, we choose three representative momenta. $K$ and $M$ lie at the edge of the Brillouin zone. $K$ is equivalent to $K^{\prime}$ by the altermagnetic symmetry. The momentum $\Gamma+\epsilon K$ is located close to the zone center slightly off in the direction of $K$.
Let us consider first the Fig. 7. The lines stand for the time-parametrized Landau-damped precession of natural magnon modes at given momentum excited at $t=0$. For a particular atomic magnetic moment, the observer is positioned against the direction of the moment in the ground state. The stars denote the initial direction of the moments at $t=0$. All three magnon modes are shown and kept apart using the color-coding from Fig. 6. Fig. 8 has the same physical content but the local $x$ and $y$ magnon components are shown as an explicit function of time.
Several features of these precession patterns point to the strong potential of KAFM as terahertz spintronic materials. The small momentum ($\Gamma+\epsilon K$) eigenvectors involve the precession of all magnetic moments. At the same time, they illustrate the pronounced chirality-dependent Landau damping in this class of systems. We clearly observe that the $\beta$-mode performs numerous revolutions prior to its decay while the $\alpha$-modes complete hardly a single one before fading away. This behavior is consistent with the FHWM discussion presented in Subsec. 4.3. Another interesting property of the magnonic eigenvectors is the magnetic sublattice localization for large momenta. At the $M$-point, one of the modes localizes on Mn1 atoms while the other two excite Mn2 and Mn3 sublattices. The phenomenon is even more pronounced at the $K$ point where in every mode only one sublattice involves in the precession. This behavior can be used to gain precise control over the dynamics of single atomic moments.
5 Summary
@@
We presented a novel implementation of the ab initio linear response time-dependent density functional theory addressing spin excitations in non-collinear magnets based on the Korringa-Kohn-Rostoker Green’s function method. We exposed the details of the numerical implementation, focusing on the managing the algorithmic complexity by means of symbolic computational algebra, and discussed the generalization of the adiabatic local spin density approximation to the exchange-correlation kernel to the non-collinear case. The emergence of multiple Nambu-Goldstone bosons in the non-collinear dynamic susceptibility calculations is elaborated on formally and from the numerical convergence point of view, being linked to the zero-energy magnetic deformation eigenvectors. With the aid of the formalism, we analyzed the spin density response of the KAFM IrMn3, in particular, the dispersion and Landau damping of magnons as well as the real-time and real-space dynamics of the spin-wave modes. Due to the non-monotonous dependence of the damping on the magnon frequency and chirality the excitations become attractive in the terahertz spintronics applications. We hope that this work will give rise to further theoretical and experimental progress in the young field of magnonics based on non-collinear ground state systems.
@@
Acknowledgments. We thank Leonid M. Sandratskii, László Szunyogh, Patrick Perndorfer, Sebastian Paischer, and Igor Maznichenko for fruitful discussions. P.B. and A.E. acknowledge the funding, respectively, by Österreichischer Fonds zur Förderung der Wissenschaftlichen Forschung (FWF) under grant I 5384 and by Deutsche Forschungsgemeinschaft (DFG) under grant BU 4062/1-1.